OpenStudy (anonymous):

How would you factor/simplify this? x^2 - 1/ x^2 + 2x +1

OpenStudy (anonymous):

x^2 - 1 = (x+1)(x-1) x^2 + 2x + 1 = (x+1)^2

OpenStudy (anonymous):

you get (x-1)/(x+1)